OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, and 3 + 9i

OpenStudy (anonymous):

note that if a polynomial has real coefficients with a complex zero \(z=a+bi\) then it is necessary that \(\bar{z}=a-bi\) is also a complex zero... \(P(a+bi)=P(a-bi)=0\) http://en.wikipedia.org/wiki/Complex_conjugate_root_theorem

OpenStudy (anonymous):

so if \(3+9i\) is a root then so is \(3-9i\) necessarily giving us:$$P(x)=(x-8)(x+14)(x-(3+9i))(x-(3-9i))$$can you distribute that out?